OpenStudy (anonymous):

Tom drove 737 miles in 11 hours At the same rate, how long would it take him to drive 603 miles?

OpenStudy (anonymous):

speed of tom=(distance/time)=(737/11) = 67 miles per hour time taken by tom=(distance/speed) =603/67= 9hour hope you got it @superduper1

OpenStudy (anonymous):

Thanks